Circular No. 6920

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

XTE J2012+381 R. Remillard, A. Levine, and A. Wood, Massachusetts Institute of

Technology (MIT), reporting for the RXTE ASM team at MIT and Goddard Space Flight Center (GSFC), write: “The RXTE All-Sky Monitor has detected a transient x-ray source at α = 20^{h}12^{m}.6, δ = +38^{o}11′ (equinox 2000.0). The position uncertainty (90-percent confidence) is ∼ ± 10′ in α and ± 4′ in δ. In daily averages, the ASM count rate (2–12 keV) was 23 mCrab on May 24.6 UT and 28 on May 25.4, rising to 55 mCrab on May 26.5 and to 88 on May 27.3. The spectrum is becoming softer as the source brightens. Observations at other wavelengths are urgently needed.”

R. M. Wagner, Ohio State University; S. Starrfield, Arizona State University; C. Shrader, GSFC; and E. Bowell, B. Skiff, and B. Koehn, Lowell Observatory, on behalf of the Lowell Observatory Near Earth Object Search (LONEOS) collaboration, report: “We obtained several unfiltered images of the RXTE/ASM error box with the LONEOS Schmidt telescope and CCD camera (field 3^{o}.2 × 1^{o}.6; scale 2′′.8/pixel) on May 27.34 UT. Comparison of these images with the Digital Sky Survey (DSS) reveals a variable stellar object and possible optical counterpart of XTE J2012+381 within the ASM error box at α = 20^{h}12^{m}30^{s}.3, δ = +38^{o}08′59′′ (equinox 2000.0; uncertainty 1′′ in both coordinates). The object appears on the DSS at mag ∼ 17–17.5 and is distinctly fainter than a star located 15′′ to the south. However, the object appears 2–3 mag brighter on the LONEOS images and comparable in brightness to the stellar object 15′′ to the south. No other obviously variable objects or new objects not present on the DSS were found in the LONEOS images of the error box.”

STAR NEAR NGC 4013 = SUPERNOVA 1989Z? R. Kushida reports via S. Nakano, Sumoto, Japan, that this star (cf.

IAUC 6911, 6914) has been present with no obvious variability on her 20 CCD frames of this galaxy taken since 1995. Y. Kushida has measured the following position of this star from a CCD frame taken with a 0.40-m f/6 reflector at Yatsugatake South Base Observatory on 1998 May 23.479 UT (V = 12.6): α = 11^{h}58^{m}32^{s}.06, δ = +43^{o}56′53′′.6 (equinox 2000.0). Images taken by the Kushidas on 1995 Jan. 10 and 1996 Jan. 3 yield V = 12.5 (and virtually identical positions). M. Aoki reports (also via Nakano) that this star is present on his CCD frames taken since 1996; on 1998 May 23.452, he finds V = 12.8 and position end figures 32^{s}.09, 53′′.6. The Kushidas indicate that this star is very near the position of SN 1989Z.

1998 May 27  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Daniel W. E. Green